TANGENTS AND NORMALS
find the points on the curve y^2=2x^3 at which the slope of the tangent is 3.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Slope = 3
y^2 = 2x^3
2yy' = 2.3x^2
2.3y = 2.3x^2
y = x^2
Put in curve
x^4 = 2x^3
x = 0, 2
y = 0, 4